OpenStudy (anonymous):

a line passes through point (-1,-6) and has a slope of -6. write an equation for this line

OpenStudy (anonymous):

because it has a slope of -6, you can use slope-intercept form to write it: y=-6x+b In this case, b is still unknown, so in order to find it, you have to plug in the x and y values of the point given (-1,-6). Thereofre: -6=-6(-1)+b -6=6+b Therefore, b=-12. So your equation would be y=-6x-12

OpenStudy (nincompoop):